# EDBench: Large-Scale Electron Density Data for Molecular Modeling

### v1

@@ 3.1 Dataset
Construction of EDBench dataset. We performed large-scale density functional theory (DFT) calculations on 3,359,472 molecules from the PCQM4Mv2 dataset (Hu et al., 2021) using Psi4 1.7 (Turney et al., 2012; Smith et al., 2020). We employed the widely used B3LYP hybrid functional due to its strong empirical performance across diverse molecular systems. The choice of reference wavefunction was determined by the spin multiplicity, which was computed from the number of unpaired electrons according to Hund’s rule. Specifically, we used a restricted Hartree-Fock (RHF) reference for closed-shell systems (multiplicity = 1), and an unrestricted Hartree-Fock (UHF) reference for open-shell systems (multiplicity > 1) to allow for independent optimization of $\alpha$ and $\beta$ spin orbitals. Basis sets were assigned based on elemental composition. We used 6-31G** for molecules without sulfur, while for sulfur-containing molecules, we used 6-31+G** to incorporate diffuse functions that better capture the more delocalized and polarizable electron distributions of heavier atoms like sulfur. After achieving self-consistent field (SCF) convergence, we generated cube files containing electron density (ED) data from Equation 1 with a grid spacing of 0.4 Bohr, a padding of 4.0 Bohr, and a density fraction threshold of 0.85 to define the isosurface region. Examples of the ED visualization are shown in Appendix A.2. All computations were carried out on a high-performance server equipped with 8 Intel(R) Xeon(R) Platinum 8270 CPUs, each with 26 physical cores and 2 threads per core, yielding a total of 416 logical cores. The total computational cost exceeded 205,000 core-hours, equivalent to approximately 23.4 years of single-core compute time. Regarding data quality and reliability, refer to Appendix A.3.
Statistical information. Figure 2(a) shows the distribution of molecular lengths in EDBench, counting only heavy atoms (excluding hydrogens), with most molecules containing no more than 20. Figure 2(b) presents the distribution of heavy atom types, covering 21 distinct elements. Figures 2(c) and 2(d) show the distributions of ED lengths and the molecular mean ED values, respectively. The average ED length exceeds the average atomic length by more than four orders of magnitude.
Figure 2: (a) Distribution of molecular lengths (heavy atoms only). (b) Distribution of heavy atom counts. (c) Distribution of ED vector lengths. (d) Distribution of per-molecule mean ED values.
@@ 3.2 Tasks
To comprehensively evaluate the capacity of the model to understand ED data, we define a suite of tasks based on both molecular structures (MS) and ED, focusing on three fundamental capabilities: prediction of quantum property, retrieval between MS and ED, and generation of ED based on MS (for the significance of these three types of tasks, see Appendix A.4). These tasks are constructed and conditionally sampled from the EDBench. To facilitate the development of ED-oriented machine learning methods within the community, we set the dataset size to a moderate scale of up to $n^{max}=50,000$ molecules, with remaining data available for future research on pre-training strategies. We use scaffold split to divide the dataset into 80% training set, 10% validation set and 10% test set, which is an out-of-distribution split setting and is widely used to evaluate the generalization ability of the model (Xiang et al., 2024a, b). We summarize the statistics of the designed datasets in Table 2 (distribution details in Appendix A.5). We next explain the construction details of these tasks.
Table 2: Statistical information of designed 6 benchmarks with a scaffold split.
@@
ED5-MM 49,917 39,933/4,992/4,992 4 Regression 4 multipole moment (3 Dipoles {X, Y, Z}, Magnitude)
@@
[…]
@@ 3.3 Methods
To assess the model’s understanding of electron density (ED), we design tailored learning paradigms for each task type (prediction, retrieval, and generation). For clarity, we formalize the molecular structure (MS) with $n$ atoms as $\mathcal{G}=(\mathcal{V},\mathcal{Z}^{\mathcal{G}})$, where $\mathcal{V}=\{v_{1},v_{2},...,v_{n}\}\in\mathbb{R}^{n\times 1}$ and $\mathcal{Z}^{\mathcal{G}}=\{z^{\mathcal{G}}_{1},z^{\mathcal{G}}_{2},...,z^{\mathcal{G}}_{n}\}\in\mathbb{R}^{n\times 3}$ denote atomic types and their corresponding 3D coordinates, respectively. The ED data with $m$ points is denoted as $\mathcal{P}=(\mathcal{Z}^{\mathcal{P}},\mathcal{D})$, where $\mathcal{Z}^{\mathcal{P}}=\{z^{\mathcal{P}}_{1},z^{\mathcal{P}}_{2},...,z^{\mathcal{P}}_{m}\}\in\mathbb{R}^{n\times 3}$ represents the ED coordinates and the corresponding density values $\mathcal{D}=\{d_{1},d_{2},..,,d_{n}\}\in\mathbb{R}^{n\times 1}$. We denote the MS encoder and ED encoder as $\texttt{Enc}_{\mathcal{G}}$ and $\texttt{Enc}_{\mathcal{P}}$, respectively, to extract latent representations from MS and ED.
@@
The models are trained with the InfoNCE loss (He et al., 2020), which pulls matched pairs closer in the embedding space while pushing apart mismatched ones. Formally, given a batch of $n$ paired samples $\{(\mathcal{G}_{i},\mathcal{P}_{i})\}_{i=1}^{n}$, the loss for a positive pair $(\mathcal{G}_{i},\mathcal{P}_{i})$ is defined as:
@@
For the generation task, we construct a heterogeneous graph (Shi et al., 2022), defined as:
@@
where $\mathcal{HG}$ contains two types of nodes: atoms and electrons. To construct the edge set $\mathcal{E}$, we perform a $k$-nearest neighbor search ($k=9$) for each node, retrieving the $k$ closest nodes of the same type and $k$ of the opposite type, which results in 18 edges per node, forming atom–atom, atom–electron, and electron–electron connections. Since the goal is to predict ED from MS, we mask all ED values to obtain the masked graph $\hat{\mathcal{HG}}$. We extend Equivariant Graph Neural Network (EGNN) (Satorras et al., 2021), called HGEGNN, to support heterogeneous graph. In HGEGNN, we treat electrons as special atoms and apply the same EGNN operations as used for regular atoms. We then input $\hat{\mathcal{HG}}$ into an HGEGNN to extract node representations $h^{\mathcal{HG}}$, which are split into atomic features $h^{\mathcal{HG}}_{\mathcal{G}}$ and electronic features $h^{\mathcal{HG}}_{\mathcal{P}}$. Finally, we apply a prediction head $\texttt{Enc}_{t}^{\text{EDP}}$ to the electronic features to generate the masked density values:
@@ 3.4 Evaluations
In the prediction tasks, the predicted labels are obtained via Equation 3. Specifically, in ED5-EC, ED5-OE, and ED5-MM, we evaluate the prediction performance using MAE between the predicted and ground-truth values, i.e., $(\hat{y}_{\mathcal{P}}^{EC},y_{\mathcal{P}}^{EC})$, $(\hat{y}_{\mathcal{P}}^{OE},y_{\mathcal{P}}^{OE})$, and $(\hat{y}_{\mathcal{P}}^{MM},y_{\mathcal{P}}^{MM})$. For ED5-OCS, we assess classification performance using accuracy, ROC-AUC, AUPR, and F1-score between the predicted logits $\hat{y}_{\mathcal{P}}^{OCS}$ and ground-truth labels $y_{\mathcal{P}}^{OCS}$. In the retrieval task, we evaluate the quality of the latent features $h_{\mathcal{G}}$ and $h_{\mathcal{P}}$ extracted via Equation 4. Specifically, in ED5-MER, given a molecular feature $h_{\mathcal{G}_{i}}$ as the anchor, we retrieve from a set of ED features by computing cosine similarities and ranking the results; Top-$k$ accuracy ($k=1,3,5$) is used as the evaluation metric, where a hit is counted if the correct match appears in the top $k$ results. Similarly, we perform retrieval in the opposite direction using $h_{\mathcal{P}_{i}}$ as the anchor and $h_{\mathcal{G}}$ as the candidate set. In the generation task, the predicted ED $\hat{\mathcal{D}}$ is obtained via Equation 7. We evaluate the generation performance using MAE, Pearson and Spearman’s rank correlation coefficients between $\hat{\mathcal{D}}$ and the ground-truth $\mathcal{D}$ in ED5-EDP.
@@ 4 Experiment and discussion
Baseline. For comprehensiveness of the evaluation, we evaluate both molecular structure-based and electron density-based methods. Specifically, we selected several state-of-the-art baselines for evaluation on the proposed benchmark: (i) two geometric models based on molecular structure (MS): GeoFormer (Wang et al., 2023a) and EquiformerV2 (Liao et al., 2024); (ii) two point cloud models based on electron density (ED): PointVector (Deng et al., 2023) and X-3D (Sun et al., 2024). GeoFormer and EquiformerV2 are Transformer-based architectures that use Interatomic Positional Encoding (IPE) and higher-degree tensors, respectively, to learn the interaction relationships between atoms. Unlike GeoFormer and EquiformerV2, which are specifically designed for molecules, PointVector and X-3D are the latest methods that focus on real-world point clouds. They are MLP(Multi-layer Perceptron)-based and explict structure-based architectures, respectively, offering excellent computational efficiency to handle large-scale point clouds. A comparison of the computational efficiency across different models is provided in Appendix A.6.
Setup. The codes of all baselines are available from their GitHub repositories and we reproduce them on our benchmarks. We use the same experimental settings as these baselines. All datasets are split using a scaffold split (Xiang et al., 2024a) based on the out-of-division (OOD) scenario, which enables evaluating the generalization of the model. We repeat the experiments three times with different random seeds and report the means and standard variances on the test set. The test set results are selected according to the best validation set performance. Due to the excessive length of the ED vectors (Figure 2(c)), we introduce a threshold $\rho_{\tau}$ to filter out electrons in regions with negligible density (all ED values below $\rho_{\tau}$ are discarded). All models were trained using either NVIDIA A100 (80GB PCIe) or GeForce RTX 3090 (24GB) GPUs, depending on their memory requirements.
@@ 4.1 Performance on prediction tasks
Table 4: The performance of MAE×$\times 100$ on 7 orbital energies of the ED5-OE with $\rho_{\tau}=0.05$.
@@
Table 5: The MAE performance on multipole moments from the ED5-MM dataset with $\rho_{\tau}=0.05$.
@@
Tables 3, 4, 5, and 6 report the performance of recent models on the ED5-EC, ED5-OE, ED5-MM, and ED5-OCS datasets, respectively. We observe that X-3D consistently outperforms PointVector, achieving the best results on ED5-EC (Table 3), ED5-MM (Table 5), and ED5-OCS (Table 6). Notably, both X-3D and PointVector exhibit significantly stronger performance on orbital energy prediction (Table 4) than on energy component prediction (Table 3). This is likely due to the stronger locality of orbital energies, which are more directly linked to local ED patterns, allowing models to extract relevant features more effectively. In contrast, predicting energy components requires integrating over the entire ED, demanding the learning of more complex global interactions. These results further validate the effectiveness of using ED as a model input and demonstrate its utility in capturing physically meaningful patterns.
@@
While X-3D and PointVector were not designed for ED data, their strong performance on our benchmarks underscores the potential of ED-based learning in quantum property prediction. We expect tailored models to further improve performance, advancing more accurate and efficient quantum modeling.
@@ 4.2 Performance on retrieval tasks
We employ {GeoFormer (G), EquiformerV2 (E)} and {PointVector (P), X-3D (X)} as the MS encoder $\texttt{Enc}_{\mathcal{G}}$ and ED encoder $\texttt{Enc}_{\mathcal{P}}$, respectively, in Equation 3. These components are paired to form four combinations: G+P, G+X, E+P, and E+X. Their retrieval performance is evaluated in Figure 3. The results show that combinations involving E (i.e., E+P and E+X) consistently outperform those involving G, highlighting the importance of selecting an appropriate encoder for retrieval tasks. Overall, the strong performance of E+P and E+X demonstrates their potential for ED-based virtual screening and MS-based electronic-level molecular understanding. For more detailed retrieval results, please refer to Appendix A.7.

### v2

@@ 3.1 Dataset
Construction of EDBench dataset. We performed large-scale density functional theory (DFT) calculations on 3,359,472 molecules from the PCQM4Mv2 dataset (hu2021ogblsc, ) using Psi4 1.7 (turney2012psi4, ; smith2020psi4, ) (We describe the reasons about why PubCheMQC (hu2021ogblsc, ) was not used in Appendix A.4). We employed the widely used B3LYP hybrid functional due to its strong empirical performance across diverse molecular systems. The choice of reference wavefunction was determined by the spin multiplicity, which was computed from the number of unpaired electrons according to Hund’s rule. Specifically, we used a restricted Hartree-Fock (RHF) reference for closed-shell systems (multiplicity = 1), and an unrestricted Hartree-Fock (UHF) reference for open-shell systems (multiplicity > 1) to allow for independent optimization of $\alpha$ and $\beta$ spin orbitals. Basis sets were assigned based on elemental composition. We used 6-31G** for molecules without sulfur, while for sulfur-containing molecules, we used 6-31+G** to incorporate diffuse functions that better capture the more delocalized and polarizable electron distributions of heavier atoms like sulfur. After achieving self-consistent field (SCF) convergence, we generated cube files containing electron density (ED) data from Equation 1 with a grid spacing of 0.4 Bohr, a padding of 4.0 Bohr, and a density fraction threshold of 0.85 to define the isosurface region. Examples of the ED visualization are shown in Appendix A.5. All computations were carried out on a high-performance server equipped with 8 Intel(R) Xeon(R) Platinum 8270 CPUs, each with 26 physical cores and 2 threads per core, yielding a total of 416 logical cores. The total computational cost exceeded 205,000 core-hours, equivalent to approximately 23.4 years of single-core compute time (See Appendix A.6 for details). Regarding data quality and reliability, refer to Appendix A.7. In addition, we verify the fidelity of different functionals to the ED in Appendix A.8 and Ti- and Zn-containing molecules in Appendix A.9.
Statistical information. Figure 2(a) shows the distribution of molecular lengths in EDBench, counting only heavy atoms (excluding hydrogens), with most molecules containing no more than 20. We explain in the Appendix A.10 why the maximum number of molecules is not more than 20. Figure 2(b) presents the distribution of heavy atom types, covering 21 distinct elements. Figures 2(c) and 2(d) show the distributions of ED lengths (the number of ED points) and the molecular mean ED values, respectively. The average ED length exceeds the average atomic length by more than four orders of magnitude.
Figure 2: (a) Distribution of molecular lengths (heavy atoms only). (b) Distribution of heavy atom counts. (c) Distribution of ED lengths. (d) Distribution of per-molecule mean ED values.
@@ 3.2 Tasks
To comprehensively evaluate the capacity of the model to understand ED data, we define a suite of tasks based on both molecular structures (MS) and ED, focusing on three fundamental capabilities: prediction of quantum property, retrieval between MS and ED, and generation of ED based on MS (for the significance of these three types of tasks, see Appendix A.11). These tasks are constructed and conditionally sampled from the EDBench. To facilitate the development of ED-oriented machine learning methods within the community, we set the dataset size to a moderate scale of up to $n^{max}=50,000$ molecules, with remaining data available for future research on pre-training strategies. We use scaffold split to divide the dataset into 80% training set, 10% validation set and 10% test set, which is an out-of-distribution split setting and is widely used to evaluate the generalization ability of the model (xiang2024image, ; xiang2024molecular, ). We summarize the statistics of the designed datasets in Table 2 (distribution details in Appendix A.12). We next explain the construction details of these tasks.
Table 2: Statistical information of designed 6 benchmarks with a scaffold split. ED5-EC, ED5-OE, ED5-MM, ED5-OCS, ED5-MER and ED5-EDP represent "Energy Components prediction", "Orbital Energy estimation", "Multipole Moment regression", "Open-/Closed-Shell classification", "MS and ED retrieval" and "ED Prediction", respectively.
@@
ED5-$M$ 49,917 39,933/4,992/4,992 4 Regression 4 multipole moment (3 Dipoles {X, Y, Z}, Magnitude)
@@
Retrieval between MS and ED. Retrieval between MS and ED is a fundamental task. Retrieving molecular structures from ED ($\texttt{ED}\rightsquigarrow\texttt{MS}$) enables electron-level virtual screening, while retrieving ED from structures ($\texttt{MS}\rightsquigarrow\texttt{ED}$) supports electron-aware models—facilitating molecular representation learning, inverse design, and quantum-informed modeling. To construct the ED5-MER dataset for bidirectional retrieval between MS and ED, we group all molecules in EDBench by structure cluster $C^{s}$ and uniformly sample $m$ anchor (MS and ED) from each group. For each anchor, we sample $n^{neg}=10$ negative samples: half from the same cluster (easy negatives) and half from different clusters (hard negatives). The final task involves identifying the correct ED (or MS) from a set of 11 candidates given an anchor MS (or ED).
Generation of ED based on MS. Generating ED from MS ($\texttt{MS}\rightarrow\texttt{ED}$) is a highly valuable task, as it can significantly reduce the computational cost associated with DFT-based ED calculations. Since ED is inherently dependent on both molecular connectivity and 3D geometry, we ensure diversity in both structure and density by grouping molecules via $C^{s}$ and uniformly sampling $m$ MS-ED pairs from each group. In this task, the model is given an MS as input and is required to predict its ED.
@@ 3.3 Methods
To assess the model’s understanding of electron density (ED), we design tailored learning paradigms for each task type (prediction, retrieval, and generation). For clarity, we formalize the molecular structure (MS) with $n$ atoms as $\mathcal{G}=(\mathcal{V},\mathcal{Z}^{\mathcal{G}})$, where $\mathcal{V}=\{v_{1},v_{2},...,v_{n}\}\in\mathbb{R}^{n\times 1}$ and $\mathcal{Z}^{\mathcal{G}}=\{z^{\mathcal{G}}_{1},z^{\mathcal{G}}_{2},...,z^{\mathcal{G}}_{n}\}\in\mathbb{R}^{n\times 3}$ denote atomic types and their corresponding 3D coordinates, respectively. The ED data with $m$ points is denoted as $\mathcal{P}=(\mathcal{Z}^{\mathcal{P}},\mathcal{D})$, where $\mathcal{Z}^{\mathcal{P}}=\{z^{\mathcal{P}}_{1},z^{\mathcal{P}}_{2},...,z^{\mathcal{P}}_{m}\}\in\mathbb{R}^{m\times 3}$ represents the ED coordinates and the corresponding density values $\mathcal{D}=\{d_{1},d_{2},..,,d_{m}\}\in\mathbb{R}^{m\times 1}$. We denote the MS encoder and ED encoder as $\texttt{Enc}_{\mathcal{G}}$ and $\texttt{Enc}_{\mathcal{P}}$, respectively, to extract latent representations from MS and ED.
@@
The models are trained with the InfoNCE loss (he2020momentum, ), which pulls matched pairs closer in the embedding space while pushing apart mismatched ones. Formally, given a batch of $n$ paired samples $\{(\mathcal{G}_{i},\mathcal{P}_{i})\}_{i=1}^{n}$, the loss for a positive pair $(\mathcal{G}_{i},\mathcal{P}_{i})$ is defined as:
@@
For the generation task, we construct a heterogeneous graph (shi2022heterogeneous, ), defined as:
@@
where $\mathcal{HG}$ contains two types of nodes: atoms and electrons. To construct the edge set $\mathcal{E}$, we perform a $k$-nearest neighbor search ($k=9$) for each node, retrieving the $k$ closest nodes of the same type and $k$ of the opposite type, which results in 18 edges per node, forming atom–atom, atom–electron, and electron–electron connections. Since the goal is to predict ED from MS, we mask all ED values to obtain the masked graph $\hat{\mathcal{HG}}$. We extend Equivariant Graph Neural Network (EGNN) (satorras2021n, ), called HGEGNN, to support heterogeneous graph. In HGEGNN, we treat electrons as special atoms and apply the same EGNN operations as used for regular atoms. We then input $\hat{\mathcal{HG}}$ into an HGEGNN to extract node representations $h^{\mathcal{HG}}$, which are split into atomic features $h^{\mathcal{HG}}_{\mathcal{G}}$ and electronic features $h^{\mathcal{HG}}_{\mathcal{P}}$. Finally, we apply a prediction head $\texttt{Enc}_{t}^{\text{EDP}}$ to the electronic features to generate the masked density values:
@@ 3.4 Evaluations
In the prediction tasks, the predicted labels are obtained via Equation 3. Specifically, in ED5-EC, ED5-OE, and ED5-MM, we evaluate the prediction performance using MAE between the predicted and ground-truth values, i.e., $(\hat{y}_{\mathcal{P}}^{EC},y_{\mathcal{P}}^{EC})$, $(\hat{y}_{\mathcal{P}}^{OE},y_{\mathcal{P}}^{OE})$, and $(\hat{y}_{\mathcal{P}}^{MM},y_{\mathcal{P}}^{MM})$. For ED5-OCS, we assess classification performance using accuracy, ROC-AUC, AUPR, and F1-score between the predicted logits $\hat{y}_{\mathcal{P}}^{OCS}$ and ground-truth labels $y_{\mathcal{P}}^{OCS}$. In the retrieval task, we evaluate the quality of the latent features $h_{\mathcal{G}}$ and $h_{\mathcal{P}}$ extracted via Equation 4. Specifically, in ED5-MER, given a molecular feature $h_{\mathcal{G}_{i}}$ as the anchor, we retrieve from a set of ED features by computing cosine similarities and ranking the results; Top-$k$ accuracy ($k=1,3,5$) is used as the evaluation metric, where a hit is counted if the correct match appears in the top $k$ results. Similarly, we perform retrieval in the opposite direction using $h_{\mathcal{P}_{i}}$ as the anchor and $h_{\mathcal{G}}$ as the candidate set. In the generation task, the predicted ED $\hat{\mathcal{D}}$ is obtained via Equation 7. We evaluate the generation performance using MAE, Pearson correlation coefficient and Spearman’s rank correlation coefficient between predicted ED $\hat{\mathcal{D}}$ and the ground-truth $\mathcal{D}$ in ED5-EDP.
@@ 4 Experiment and discussion
Baseline. For comprehensiveness of the evaluation, we evaluate both molecular structure-based and electron density-based methods. Specifically, we selected several state-of-the-art baselines for evaluation on the proposed benchmark: (i) two geometric models based on molecular structure (MS): GeoFormer (wang2023geometric, ) and EquiformerV2 (liao2024equiformerv, ); (ii) two point cloud models based on electron density (ED): PointVector (Deng_2023_CVPR, ) and X-3D (sun2024x, ). GeoFormer and EquiformerV2 are Transformer-based architectures that use Interatomic Positional Encoding (IPE) and higher-degree tensors, respectively, to learn the interaction relationships between atoms. Unlike GeoFormer and EquiformerV2, which are specifically designed for molecules, PointVector and X-3D are the latest methods that focus on real-world point clouds. They are MLP(Multi-layer Perceptron)-based and explict structure-based architectures, respectively, offering excellent computational efficiency to handle large-scale point clouds. A comparison of the computational efficiency across different models is provided in Appendix A.13.
Setup. The codes of all baselines are available from their GitHub repositories and we reproduce them on our benchmarks. We use the same experimental settings as these baselines. All datasets are split using a scaffold split (xiang2024image, ) based on the out-of-division (OOD) scenario, which enables evaluating the generalization of the model. We repeat the experiments three times with different random seeds and report the means and standard variances on the test set. The test set results are selected according to the best validation set performance. Due to the excessive length of the ED vectors (Figure 2(c)), we introduce a threshold $\rho_{\tau}$ to filter out electrons in regions with negligible density (all ED values below $\rho_{\tau}$ are discarded). All models were trained using either NVIDIA A100 (80GB PCIe) or GeForce RTX 3090 (24GB) GPUs, depending on their memory requirements.
@@ 4.1 Performance on prediction tasks
[…]
@@
Table 5: The MAE performance on multipole moments from the ED5-$M$ dataset with $\rho_{\tau}=0.05$.
@@
Tables 3, 4, 5, and 6 report the performance of recent models on the ED5-EC, ED5-OE, ED5-MM, and ED5-OCS datasets, respectively. We observe that X-3D consistently outperforms PointVector, achieving the best results on ED5-EC (Table 3), ED5-MM (Table 5), and ED5-OCS (Table 6). Notably, both X-3D and PointVector exhibit significantly stronger performance on orbital energy prediction (Table 4) than on energy component prediction (Table 3). This is likely due to the stronger locality of orbital energies, which are more directly linked to local ED patterns, allowing models to extract relevant features more effectively. In contrast, predicting energy components requires integrating over the entire ED, demanding the learning of more complex global interactions. In addition, see Appendix A.14 for the results of a training with full corpus (about 2.67 million molecules) on ED5-OE task and see Appendix A.15 for extension effectiveness study to periodic systems (materials). Overall, these results validate the effectiveness of using ED as a model input and demonstrate its utility in capturing physically meaningful patterns.
@@ 4.2 Performance on retrieval tasks
We employ {GeoFormer (G), EquiformerV2 (E)} and {PointVector (P), X-3D (X)} as the MS encoder $\texttt{Enc}_{\mathcal{G}}$ and ED encoder $\texttt{Enc}_{\mathcal{P}}$, respectively, in Equation 3. These components are paired to form four combinations: G+P, G+X, E+P, and E+X. Their retrieval performance is evaluated in Figure 3. The results show that combinations involving E (i.e., E+P and E+X) consistently outperform those involving G, highlighting the importance of selecting an appropriate encoder for retrieval tasks. Overall, the strong performance of E+P and E+X demonstrates their potential for ED-based virtual screening and MS-based electronic-level molecular understanding. For more detailed retrieval results, please refer to Appendix A.16 and see Appendix A.17 for the sensitivity analysis of negative samples.